Find the (exact) value of the six trigonometric ratios associated with the angle described. The terminal side of an angle contains the point (-3,4).
Answer by lwsshak3(11628)

Find the (exact) value of the six trigonometric ratios associated with the angle described. The terminal side of an angle contains the point (-3,4)
**
you are working with a 3-4-5 right triangle in quadrant II where sin and csc>0, and cos, sec, tan, and cot<0
..
sin=4/5
cos=-3/5
tan=-4/3
csc=5/4
sec=-3/5
cot=-3/4